OpenStudy (anonymous):

An object travels with velocity v = 4.0 meters/second and it makes an angle of 60.0° with the positive direction of the y-axis. Calculate the possible values of vx.

OpenStudy (anonymous):

\[v_x=v \cos \theta=4.0 \cos 60^0=2.0 \text{ m/s}\]